# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v5

@@ IV-E1 SlideMatch
The SlideMatch algorithm takes a pair of object-level metric-semantic maps from either the same robot (i.e., historical map and current map for loop closure) or a pair of different robots. Similar to our previous work in [63], SlideMatch algorithm checks within a search region if a valid loop closure is found and outputs the relative transformation between the reference frames of these two maps. Unlike our previous work [63], here we also leverage additional information such as semantic labels and dimensions of objects in the maps. The search region of the algorithm encompasses continuous ranges of X and Y positions and yaw angles, where the relative transformation between the two maps may exist. The search procedure is carried out over the search region using a user-defined search resolution for discretization.
@@
Preprocessing: For inter-robot place recognition, we designed a preprocessing step that zero-centers the two object-level maps before performing data association. This step reduces the search region and computation of the SlideMatch algorithm, especially when the two robots start far apart. Furthermore, it enables the algorithm to automatically calculate the search region by setting the X and Y search ranges to encompass the region where the two maps overlap and the yaw search range to $(-\pi,+\pi]$.
@@
Anytime implementation: The anytime implementation progressively improves the quality of place recognition until the user-specified compute budget is exhausted. Quality is measured by the number of inlier matches. This approach guarantees the algorithm returns the best estimate within the given computation time budget. A valid loop closure is considered to be found if the matching score of the best estimate exceeds the valid loop closure inlier threshold, which will be detailed below. This anytime procedure is achieved by gradually increasing the search region for data association until it covers the entire region or the compute budget is used up.
@@
Data association: The data association step involves finding the possible set of landmarks that are common in both maps. Candidate transformation samples are generated from the search region by discretizing the search space across the X, Y, and yaw dimensions, based on the user-specified resolutions. Specifically, the algorithm iterates through all possible X and Y position and yaw angle samples within the current region calculated by the anytime process. Within each iteration, the objects in the first map is transformed using the sampled X, Y, and yaw, and matched against the objects in the second map. The algorithm then evaluates the matching score by summing up the number of valid object matches. Similar to the data association in Section IV-D, a valid object model match must have the same semantic label, and their centroid distance and model difference are within the given threshold. At the end of each iteration, if the matching score exceeds the current best matching score, it will be recorded along with the valid object matches. After all iterations, the best matching score will be compared with the valid loop closure threshold to determine whether a valid loop closure has been identified.
@@ IV-E2 SlideGraph
The SlideGraph algorithm, unlike the exhaustive search-based SlideMatch algorithm, adopts a descriptor-based matching approach and combines it with a robust outlier rejection method, CLIPPER [32], for efficient and robust loop closures. While the CLIPPER framework can achieve robust data association by formulating the problem as a maximum clique search in a weighted graph. When an initial data association is unavailable, CLIPPER can still operate under an all-to-all association assumption. However, this quickly becomes too memory demanding and computationally expensive as the number of object landmarks increases in the map.
To address these limitations, we propose a method for generating initial data associations using descriptor matching. Inspired by prior research [36], our descriptor design applies Delaunay triangulation to two sets of object landmarks derived from metric-semantic maps. Candidate simplices, either triangles (2D landmarks) or tetrahedrons (3D landmarks), from each set are compared based on their sorted vertex-to-centroid distances. A valid match is found if the distance metrics between two simplices fall within a predefined threshold. From these matched simplices, the corresponding vertices are extracted as an initial set of object landmark associations. The semantic class labels of the object landmarks are used to further reject false candidate matches. This initial matching step significantly reduces the search space for the CLIPPER framework, which subsequently refines these matches by filtering out false associations through its graph-theoretic optimization. This two-stage approach ensures robust data association while maintaining computational feasibility, even for large landmark sets.
@@ IV-E3 Least square optimization
Table III: Robot platform details. Specifications on three types of robot platforms as shown in Fig. 1.
@@ V-B Deployment for autonomous exploration
We integrate the metric-semantic SLAM algorithm with our autonomous exploration and navigation systems for all three types of robot platforms. The integration process and the use of semantic maps differ for different sensing modalities. We refer the reader to our previous work [3] for active metric-semantic SLAM with LiDAR-based UAVs. The RGBD-based UAV and UGV platforms, due to larger drifts in odometry, use the outputs of the metric-semantic SLAM framework to not only explore frontiers to gather more information but also actively establish semantic loop closures so that the state estimation uncertainty can be reduced. We refer the reader to our previous work for more details [1]. Building upon this foundation, we extend [1] by (a) upgrading the SLAM module with the metric-semantic SLAM framework proposed in this work so that multi-robot collaboration can be achieved, (b) deploying the active metric-semantic SLAM system that was originally designed for Falcon 250 UAV on a UGV platform (i.e., the Scarab UGV platform), and (c) supporting a heterogeneous team of aerial and ground robots to share information and collaboratively construct metric-semantic maps. An illustration of metric-semantic SLAM-in-the-loop autonomous exploration can be found in Fig. 5.
@@ V-C4 Experiment 4
The urban environment experiments were conducted at the University of Pennsylvania’s Pennovation campus in Pennsylvania, USA. Forest experiments were conducted in the Wharton State Forest in New Jersey, USA.
V-D Evaluation metrics
The following metrics are chosen to evaluate the performance of our system: (1) Inter-robot localization accuracy. (2) Precision, recall, and F-1 score for object mapping. (3) Runtime and communication bandwidth requirements. In the rest of this section, we present a detailed analysis of the results for each experiment.
@@ V-D Evaluation metrics
Table IV: Inter-robot localization errors. SM and SG refer to the SlideMatch and SlideGraph algorithms, respectively. Position error is calculated as the L2 norm of the errors in the X, Y, and Z coordinates of the relative transformation between two robots’ reference frames. In general, the SlideGraph algorithm is more efficient and requires less parameter tuning compared to SlideMatch, while still providing comparable results. However, both algorithms encounter difficulties in scenarios where object landmarks are noisy, particularly with RGBD-based sensing. SlideGraph is relatively more prone to failures in such conditions due to insufficient pairwise consistency among object landmarks.
V-E Indoor-outdoor LiDAR-RGBD metric-semantic SLAM without inter-robot loop closures
@@
Fig. 2 shows qualitative results on metric-semantic SLAM for experiment 1.a as described in Section V-C. Outdoor objects such as cars, tree trunks, and light poles and indoor objects such as chairs, tables, and monitors are detected and mapped using our sparse metric-semantic map representation. Object models for vehicles are represented by blue cuboids, while tree trunks and light pole landmarks are shown as pink cylinders. Chairs, tables, and monitors are initially mapped using an ellipsoid model and subsequently replaced with a Computer-Aided Design (CAD) model. These models are colored blue, green, and red, respectively, according to their detected semantic class. This experiment shows the capabilities of our metric-semantic SLAM framework in integrating data from multiple robots equipped with either LiDARs or RGBD cameras, enabling the construction of a unified metric-semantic map that includes both indoor and outdoor areas.
V-F Urban and forest LiDAR-only metric-semantic SLAM with inter-robot loop closures
@@
Figure 10: Metric semantic map constructed in real time onboard three Scarab UGVs. The UGVs autonomously explore the environment while communicating with each other. Blue, orange and red trajectories correspond to three trajectories from the three UGVs. Our semantic place recognition algorithm was able to perform inter-robot loop closure, which is used to construct a merged metric-semantic map. The gray-colored accumulated point cloud is not generated by these UGVs but by a LiDAR robot for visualization purposes. The bottom panel overlays the trajectories with the metric-semantic map to qualitatively assess the validity of the results.
@@
V-G Indoor RGBD-only aerial-ground metric-semantic SLAM with inter-robot loop closures
@@
V-H Indoor LiDAR-RGBD metric-semantic SLAM with inter-robot loop closures
@@
V-I Object mapping quantitative results
@@
V-J Semantic KITTI results
In addition to conducting experiments with our own robots, we also tested our algorithm framework on the publicly available Semantic KITTI dataset. The framework takes in the point clouds, LiDAR odometry, and ground truth segmentation, and estimates the metric-semantic map as well as the robot trajectory.
The qualitative results, as shown in Fig. 7, demonstrate the ability of the proposed framework to construct metric-semantic maps and merge measurements from multiple robots. We conducted a quantitative evaluation of our SlideGraph algorithm for inter-robot loop closure using sequences #$\#5$ and #$\#7$. Sequence #$\#7$ was divided into three intervals, i.e., (0, 600), (400, 1000), and (500, 1100), for the first, second, and third robots, respectively. Sequence #$\#5$ was similarly divided into three intervals, i.e., (0, 1000), (700, 2000), and (1600, 2760), to create datasets for each of the three robots. Our algorithm successfully identified inter-robot loop closures and accurately estimated the relative transformations between the robots. For sequence #$\#5$, the SlideGraph algorithm results in a position error of 0.03 m and 0.05 m, and a rotation error of 0.080 and 0.065 for the transformations between robot 0 and robot 1, and between robot 1 and robot 2, respectively. In sequence #$\#7$, the algorithm results in a position error of 0.093 m and 0.095 m, and a rotation error of 0.096 and 0.097.
V-K Communication usage and runtime analysis
V-K1 Discussion on scalability
@@ V-K2 Experimental results
Table VI: Runtime and communication results. The number of objects is the count of object landmarks in our metric-semantic map. The runtime for factor addition and graph optimization accounts for handling both the observation of new landmarks and loop closure scenarios. The runtime on semantics-driven place recognition and loop closure algorithms includes both loop closure detection and relative transformation estimation. Note that the place recognition and loop closure process runs on a separate thread. Therefore, its delay does not impact real-time operation. SlideGraph requires significantly lower computational runtime, particularly in environments with fewer object landmarks. This efficiency is attributed to the use of descriptor-based matching, where computational demands are influenced not by the scale of the environment but only by the number of object landmarks. In contrast, SlideMatch exhaustively searches across the entire search region, and its runtime depends on various factors, such as the search discretization. The communication usage is measured by the size of messages published by each robot. Here, the average usage is calculated by dividing the total size of all messages by the number of messages, while the maximum size corresponds to the size of the last messages since they usually contain the most observations.
Additionally, the runtime for each module, including object detection and modeling, factor adding and graph optimization, and loop closure (if applicable), is reported in Table VI. Except for the loop closure module, which is used to identify inter-robot loop closures and operates on a separate CPU thread, all other modules are triggered each time new observations arrive.
The front-end runtime depends highly on the object segmentation and detection model. Our results show that on a single thread of Intel i7-10750H CPU, the Rangenet++ [56] front-end with our custom darknet-13 model runs at 1.5 hertz (Hz). The front-end with the YOLOv8 [58], specifically, YOLOv8m, runs at 1 Hz per CPU thread. When multiple CPU threads are used, the inference speed can be increased. For example, YOLOv8m can run at 3 Hz using 4 out of 12 CPU threads.
The back-end factor graph adding and graph optimization processes show different patterns across different environments. For the indoor environment, the average runtime of graph optimization is 2.15 ms, with 60 landmarks, with an average time per landmark of 0.036 ms.
The outdoor urban environment exhibits an average runtime of graph optimization of 10.14 ms, with 249 landmarks. The average time per landmark in this setting is 0.041 ms. Due to the increased density of landmarks in the outdoor forest environment, the average runtime spikes to 36.46 ms, with a significantly higher number of landmarks at 860, and an average time per landmark of 0.042 ms.
@@
Although the runtime of the semantics-driven place recognition and loop closure algorithms is longer than that of factor graph optimization, the resulting delay is acceptable. In addition, since the semantics-driven place recognition and loop closure algorithm operate in a separate CPU thread, this delay does not directly impact real-time localization and navigation. Once the computation is complete, the estimated loop closure information can be used by the factor graph to update both the map and the robot trajectories.
@@
V-L Analysis of SlideMatch and SlideGraph algorithms
@@
VI Conclusion

### v6

@@ IV-E1 SlideMatch
The SlideMatch algorithm takes a pair of object-level metric-semantic maps from either the same robot (i.e., historical map and current map for loop closure) or a pair of different robots. SlideMatch algorithm checks within a search region if a valid loop closure is found and outputs the relative transformation between the reference frames of these two maps. Unlike [67], here we also leverage additional information such as semantic labels and dimensions of objects in the maps. The search region of the algorithm encompasses continuous ranges of X and Y positions and yaw angles, where the relative transformation between the two maps may exist. The search procedure is carried out over the search region using a user-defined search resolution for discretization.
@@
Preprocessing: For inter-robot place recognition, we designed a preprocessing step that zero-centers the two object-level maps before performing data association. This step reduces the search region and computation of the SlideMatch algorithm, especially when the two robots start far apart. Furthermore, it enables the algorithm to automatically calculate the search region by setting the X and Y search ranges to encompass the region where the two maps overlap and the yaw search range to $(-\pi,\pi]$.
@@
Anytime implementation: The anytime implementation progressively improves the quality of place recognition until the user-specified compute budget is exhausted. Quality is measured by the number of inlier matches. This approach guarantees the algorithm returns the best estimate within the given computation time budget. A valid loop closure is considered to be found if the matching score of the best estimate exceeds the valid loop closure inlier threshold, which is described in more detail below. This anytime procedure is achieved by gradually increasing the search region for data association until it covers the entire region or the compute budget is used up.
@@
Data association: The data association step involves finding the possible set of landmarks that are common in both maps. Candidate transformation samples are generated from the search region by discretizing the search space across the X, Y, and yaw dimensions, based on the user-specified resolutions. Specifically, the algorithm iterates through all possible X and Y position and yaw angle samples within the current region calculated by the anytime process. Within each iteration, the objects in the first map is transformed using the sampled X, Y, and yaw, and matched against the objects in the second map. The algorithm then evaluates the matching score, which is the total sum of all valid object matches between the two maps. Similar to the data association in Section IV-D, a valid object model match must have the same semantic label, and their centroid distance and model difference are within the given threshold. At the end of each iteration, if the matching score exceeds the current best matching score, it will be recorded along with the valid object matches. After all iterations, the best matching score will be compared with the valid loop closure threshold to determine whether a valid loop closure has been identified.
@@ IV-E2 SlideGraph
The SlideGraph algorithm, unlike the exhaustive search-based SlideMatch algorithm, adopts a descriptor-based matching approach and combines it with a robust outlier rejection method, CLIPPER [36], for efficient and robust loop closures. While the CLIPPER framework can achieve robust data association by formulating the problem as a maximum clique search in a weighted graph, when an initial data association is unavailable, CLIPPER can still operate under an all-to-all association assumption. However, this quickly becomes too memory demanding and computationally expensive as the number of object landmarks increases in the map.
To address these limitations, we propose a method for generating initial data associations using descriptor matching. Inspired by prior research [42], our descriptor design applies Delaunay triangulation to two sets of object landmarks derived from metric-semantic maps. Candidate simplices, either triangles (2D landmarks) or tetrahedrons (3D landmarks), from each set are compared based on their sorted vertex-to-centroid distances. A valid match is found if the distance metrics between two simplices fall within a predefined threshold. From these matched simplices, the corresponding vertices are extracted as an initial set of object landmark associations. The semantic class labels of the object landmarks are used to further reject false candidate matches. This initial matching step significantly reduces the search space for the CLIPPER framework, which subsequently refines these matches by filtering out false associations through its graph-theoretic optimization. The total number of valid matches resulting from this refinement is treated as the matching score, which is then compared against a threshold to determine whether a valid loop closure has been detected. This two-stage approach ensures robust data association while maintaining computational feasibility, even for large landmark sets.
@@ IV-E3 Least square optimization
Table III: Robot platform details. Specifications on three types of robot platforms used in our experiments.
@@ V-B Deployment for autonomous exploration
We integrate the metric-semantic SLAM algorithm with our autonomous exploration and navigation systems for all three types of robot platforms. The integration process and the use of semantic maps differ for different sensing modalities. In [3], an active metric-semantic SLAM system with similar map representations is proposed for LiDAR-based UAVs. In contrast, the RGB-D-based UAV and UGV platforms in [1] face larger odometry drifts. To adress this, they use the outputs of the metric-semantic SLAM framework to not only explore frontiers to gather more information but also actively establish semantic loop closures so that the state estimation uncertainty can be reduced. We extend [1] by (a) upgrading the SLAM module with the metric-semantic SLAM framework proposed in this work so that multi-robot collaboration can be achieved, (b) deploying the active metric-semantic SLAM system that was originally designed for Falcon 250 UAV on a UGV platform (i.e., the Scarab UGV platform), and (c) supporting a heterogeneous team of aerial and ground robots to share information and collaboratively construct metric-semantic maps. An illustration of metric-semantic SLAM-in-the-loop autonomous exploration can be found in Fig. 5.
@@ V-C4 Experiment 4
Table IV: Inter-robot localization errors. SM and SG refer to the SlideMatch and SlideGraph algorithms, respectively. Position error is calculated as the L2 norm of the errors in the X, Y, and Z coordinates of the relative transformation between two robots’ reference frames. In general, the SlideGraph algorithm is more efficient and requires less parameter tuning compared to SlideMatch, while still providing comparable results. However, both algorithms encounter difficulties in scenarios where object landmarks are noisy, particularly with RGBD-based sensing. SlideGraph is relatively more prone to failures in such conditions due to insufficient pairwise consistency in the positions of object landmarks.
@@
V-D Indoor-outdoor LiDAR-RGBD metric-semantic SLAM without inter-robot loop closures
@@
Fig. 2 shows qualitative results on metric-semantic SLAM for experiment 1.a as described in Section V-C. Outdoor objects such as cars, tree trunks, and light poles and indoor objects such as chairs, tables, and monitors are detected and mapped using our sparse metric-semantic map representation. Object models for vehicles are represented by blue cuboids, while tree trunks and light pole landmarks are shown as pink cylinders. The ellipsoid models of chairs, tables, and monitors are replaced with corresponding Computer-Aided Design (CAD) models. This is done only for visualization purposes and an intuitive interpretation of the metric-semantic map, especially in highly cluttered indoor environments. These models are colored blue, green, and red, respectively, according to their detected semantic class. This experiment shows the capabilities of our metric-semantic SLAM framework in integrating data from multiple robots equipped with either LiDARs or RGBD cameras, enabling the construction of a unified metric-semantic map that includes both indoor and outdoor areas.
V-E Urban and forest LiDAR-only metric-semantic SLAM with inter-robot loop closures
@@
Figure 10: Metric semantic map constructed in real time onboard three Scarab UGVs. The UGVs autonomously explore the environment while communicating with each other. Blue, orange and red trajectories correspond to three trajectories from the three UGVs. Our semantic place recognition algorithm was able to perform inter-robot loop closure, which is used to construct a merged metric-semantic map. The gray-colored accumulated point cloud is not generated by these UGVs but by a LiDAR robot for visualization purposes. The figure overlays the trajectories with the metric-semantic map to qualitatively assess the validity of the results.
@@
V-F Indoor RGBD-only aerial-ground metric-semantic SLAM with inter-robot loop closures
@@
V-G Indoor LiDAR-RGBD metric-semantic SLAM with inter-robot loop closures
@@
V-H Object mapping quantitative results
@@
V-I Communication usage and runtime analysis
V-I1 Discussion on scalability
@@ V-I2 Experimental results
Table VI: Runtime and communication results. The runtime for factor addition and graph optimization accounts for handling both the observation of new landmarks and loop closure scenarios. The runtime of semantics-driven place recognition and loop-closure algorithms includes both loop closure detection and relative transformation estimation. Note that the place-recognition and loop-closure process runs on a separate thread. Therefore, its latency does not impact real-time operation. SlideGraph requires significantly lower computational runtime, particularly in environments with fewer object landmarks. This efficiency is attributed to the use of descriptor-based matching, where computational demands are influenced not by the scale of the environment but only by the number of object landmarks. In contrast, SlideMatch exhaustively searches across the entire search region, and its runtime depends on various factors, such as the search discretization. The communication usage is measured by the size of messages published by each robot. Here, the average usage is calculated by dividing the total size of all messages by the number of messages, while the maximum size corresponds to the size of the last messages since they usually contain the most observations.
@@
Additionally, the runtime for factor adding and graph optimization, and loop closure (if applicable), is reported in Table VI. Except for the loop closure module, which is used to identify inter-robot loop closures and operates on a separate CPU thread, all other modules are triggered each time new observations arrive.
The front-end runtime depends on the object segmentation and detection model. Our results show that on a single thread of Intel i7-10750H CPU, the Rangenet++ [62] front-end with our custom darknet-13 model runs at 1.5 hertz (Hz). The front-end with the YOLOv8 [64], specifically, YOLOv8m, runs at 1 Hz per CPU thread. When multiple CPU threads are used, the inference speed can be increased. For example, YOLOv8m can run at 3 Hz using 4 out of 12 CPU threads. The cuboid, cylinder, and ellipsoid modeling process runs at the corresponding object detection rate.
The back-end factor graph adding and graph optimization processes show different patterns across different environments. For the indoor environment, the average runtime of graph optimization is 2$\sim$4 ms, with 68 landmarks. The outdoor urban environment exhibits an average runtime of graph optimization of 10.14 ms, with 249 landmarks. The average time per landmark in this setting is 0.041 ms. Due to the increased density of landmarks in the outdoor forest environment, the average runtime spikes to 36.46 ms, with a significantly higher number of landmarks at 860.
@@
Although the runtime of the semantics-driven place recognition and loop closure algorithms is longer than that of factor graph optimization, the resulting latency is acceptable. In addition, since the semantics-driven place recognition and loop closure algorithm operate in a separate CPU thread, this latency does not directly impact real-time localization and navigation. Once the computation is complete, the estimated loop closure information can be used by the factor graph to update both the map and the robot trajectories.
@@
V-J Analysis of SlideMatch and SlideGraph algorithms
@@
V-K Benchmarks
In this section, we present a series of benchmark experiments evaluating our system’s performance in terms of localization accuracy, computational efficiency, and memory usage. We report results from (1) evaluating our framework on standard public datasets and (2) comparing it with other open-source methods. We note that benchmarking the entire system is challenging due to the differences in components across other metric-semantic SLAM systems. Therefore, we opt to isolate and benchmark key modules instead. This allows for a fair and focused comparison of our system.
Table VII: Benchmark comparison between SlideSLAM (SlideGraph) and existing methods. For these benchmarks, we report results between robot 0 and robot 1, as referenced by the Semantic KITTI experiments in Section V-K1 and forest experiments in Section V-E. Note that the map sizes for both CLIPPER and SlideSLAM are the same, as they use the same object maps as inputs. Also, note that for the memory and runtime benchmarks of LCDNet, the results reported are the memory consumed for each individual point cloud and the runtime to process a pair of point clouds. In practice, it scales linearly with the number of point clouds. Cells filled with a “-” indicate that the algorithm failed.
Metric Trans. Err (m) Abs. Yaw. Err (degree) Map / Input Size Runtime (ms)
Method CLIPPER LCDNet Ours CLIPPER LCDNet Ours CLIPPER (full maps) LCDNet (per scan) Ours (full maps) CLIPPER (CPU) LCDNet (per scan pair on GPU) Ours (CPU)
KITTI 05 0.04 0.09 0.03 0.04 0.22 0.08 7.2 KB 0.84 MB 7.2 KB 152 1580 40
KITTI 07 0.06 0.05 0.09 0.01 0.21 0.09 7.3 KB 0.84 MB 7.3 KB 202 1580 50
Forest - 1.10 0.13 - 0.1 2.6 16 KB 0.84 MB 16 KB - 1580 209
V-K1 Bechmark on Semantic KITTI
In addition to conducting experiments with our own robots, we also tested our algorithm framework on the publicly available Semantic KITTI dataset. The framework takes in the point clouds, initial pose estimation from SuMa++ [24], and ground truth segmentation, and estimates the metric-semantic map as well as the robot trajectory.
The qualitative results, as shown in Fig. 6, demonstrate the ability of the proposed framework to construct metric-semantic maps and merge measurements from multiple robots. We conducted a quantitative evaluation of our SlideGraph algorithm for inter-robot loop closure using sequences #$\#5$ and #$\#7$. Sequence #$\#7$ was divided into three intervals, i.e., (0, 600), (400, 1000), and (500, 1100), for the first, second, and third robots, respectively. Sequence #$\#5$ was similarly divided into three intervals, i.e., (0, 1000), (700, 2000), and (1600, 2760), to create datasets for each of the three robots. Our algorithm successfully identified inter-robot loop closures and accurately estimated the relative transformations between the robots. For sequence #$\#5$, the SlideGraph algorithm results in a position error of 0.03 m and 0.05 m, and a rotation error of 0.080 and 0.065 for the transformations between robot 0 and robot 1, and between robot 1 and robot 2, respectively. In sequence #$\#7$, the algorithm results in a position error of 0.093 m and 0.095 m, and a rotation error of 0.096 and 0.097. For sequence #5 and sequence #7, the total trajectory lengths of three robots are 2.8 km and 1.2 km. The absolute trajectory errors of the trajectories of all three robots are 0.998 m (vs. 1.154 m from SuMa++) and 0.962 m (vs. 1.148 m from SuMa++), respectively. The relative translation errors are 0.111% and 0.255%. The relative rotation errors are 0.0011 degrees/m and 0.00256 degrees/m.
V-K2 Benchmark against existing loop closure methods
To provide a comprehensive comparison, we benchmark SlideSLAM against two existing approaches from the literature. These include: (1) LCDNet [39], a deep-learning-based loop closure method for LiDAR data; and (2) CLIPPER [36], a graph-theoretic algorithm for data association.
LCDNet performs scan-to-scan place recognition using a learned 3D feature extractor and registration network. We benchmark the algorithm using weights that are pretrained on the KITTI dataset. We use the same sequences and splits of Semantic KITTI as described in the previous benchmark. We also benchmark LCDNet on our forest data, as described in the previous sections. The benchmarks for LCDNet were conducted on a GPU-equipped desktop (with an RTX 3070 GPU). The inter-robot localization results are shown in Table VII. The results are comparable in the KITTI environment. However, for the forest environment, SlideSLAM provides better localization results. This shows the advantage of our method to generalize to different environments as long as the objects are detected. Our method requires no additional fine‑tuning data for the loop closure step, whereas LCDNet does. Moreover, we also benchmark the memory requirements to store the observations or maps required for the loop closure step. For LCDNet, these representations can be individual processed point cloud scans (0.84 MB), point features extracted from the PV-RCNN network for each point cloud scan (4.18 MB), or the final global descriptors of those features for each point cloud scan (10.0 MB). As a result, enabling inter-robot loop closure detection with LCDNet requires a minimum bandwidth of 0.84 MB per LiDAR scan. In contrast, as shown in the table, the entire object map of SlideSLAM requires less than 10 KB for the KITTI sequences, which contain hundreds of LiDAR scans. The sparsity of our map representation offers a key advantage for large-scale multi-robot experiments, as it significantly reduces both computational and communication requirements compared to methods operating on less compact map representations.
CLIPPER [36] serves as the data association backbone of our SlideSLAM (SlideGraph) module. However, the original CLIPPER lacks semantic information and relies on exhaustive pairwise matching over all object pairs, resulting in high computational demands in complex scenes (e.g., each of our forest maps consists of hundreds of landmarks). Our SlideGraph enhances CLIPPER by using semantic class labels and polygon-descriptor-based initial data association, enabling us to significantly reduce the hypothesis space. CLIPPER is benchmarked on similar sequences on Semantic KITTI as detailed in the previous section, and also on our forest dataset. The results are shown in Table VII. In challenging environments such as forests, the original CLIPPER algorithm encountered memory overflows due to the large number of landmarks, which significantly expanded the hypothesis space. Take the forest environment as an example: a pair of maps with approximately 300 and 800 landmarks would generate around 240,000 hypotheses. In contrast, our SlideGraph algorithm prunes this raw hypothesis space down to approximately 3,000, significantly reducing the size of the consistency graph by 2 orders of magnitude. As a result, SlideGraph consistently achieved robust performance with lower memory usage and shorter runtimes.
VI Discussion on Limitations and Future Work
In this section, we address a few limitations of the current system and how they motivate our future work. Firstly, the proposed sparse map representation is limited to only three simple shapes for modeling objects. While it helps maintain sparsity, the map may not always be visually informative with these shapes, and they may fail to model the finer metric details for objects with complex shapes, which are needed for downstream tasks such as robot manipulation. In addition, they may lead to information loss that results in inaccuracies or failures to identify loop closures, especially in environments with ambiguous or repetitive object layouts. Semantic aliasing may also occur when different objects are represented by the same primitive, increasing dependence on accurate semantic labeling. Misclassifications in such cases can cause errors. We are interested in investigating better map representations that can model generic objects in detail while still maintaining relative sparsity. Secondly, the current inter-robot loop closure strategy does not follow a conventional “filtering” approach, where multiple loop closures are first established and then a consensus is formed to choose the best loop closure result and filter out false positives. Currently, we only use a sufficient inlier threshold to filter out false positives and establish loop closure in a “one-shot” manner. This may not always be effective in certain situations where the environment is cluttered with the same semantic landmarks and similar pairwise distances (e.g., two conference rooms with similar table and chair arrangements, or dense forests). Therefore, we would like to employ a filtering-based approach to our current loop closure strategy, where the best loop closure candidate is chosen based on a maximum consensus of potential loop closure options. Lastly, our current place recognition and mapping formulation doesn’t account for dynamic objects. Such objects (e.g., a car changing its parking spot over time) violate the pairwise distance consistency assumption required for robust place recognition and can therefore induce faulty loop closures. In the future, we would like to leverage semantic knowledge of objects, combined with intelligent agents such as Large Language Models (LLMs), to better handle dynamic objects. Using the semantic knowledge, an LLM can infer which objects can be dynamic in nature and automatically filter them out during place recognition.
VII Conclusion